# Minimal Interaction Edge Tuning: A New Paradigm for Visual Adaptation

### v2

@@ 4.1 Experiments on VTAB-1K
We compare MIET with various fine-tuning methods. Basically, we divide them into two categories: those potentially support edge tuning and those do not. For the first category, we have linear probing, Partial-1 (which unfreeze the final layer of ViT-B), LST [27], LAST [29] and our method MIET. For the second category, we have BitFit [33], VPT [15], LoRA [14], AdaptFormer [4], FacT [16], SPT-LoRA [8].
@@ Results.
The results on VTAB-1K is shown in Table 1. Methods that support edge tuning are marked with $\checkmark$ in corresponding cell. In the “Transfer Overhead” column, we list the communication cost per image between cloud and edge devices. MIET only requires to transfer 0.58MB of information per image (which roughly equals the size of an original image), lower than other compared methods (except linear probing). This overhead is only about $1/7$ of LAST and $1/13$ of LST. Hence, MIET has met our “minimal interaction” goal.
With respect to trainable parameters, MIET only introduces 0.38M trainable parameters, lower than other methods that support edge tuning. Though MIET requires slightly more trainable parameters than some non-edge tuning methods, this difference is in fact negligible with respect to parameter size of the pretrained model (e.g., MIET:0.38/85.8=0.44% vs. FacT:0.07/85.8=0.08%, both $<1\%$).
We also list the MACs and GPU memory footprint for different methods. For non-edge tuning methods, the whole model are enforced to be fine-tuned on edge devices. For methods that support edge tuning, pretrained model is deployed on a cloud server and edge networks are deployed on edge devices, leading to very low FLOPs and memory footprint on edge devices. MIET requires the lowest MACs and GPU memory (again except linear probing) within edge devices, which are respectively about 1% and 10% of normal PEFT methods.
MIET achieves the highest group-wise accuracy compared to other methods. It has the highest group accuracy for specialized datasets and structured datasets. This result proves that MIET has strong adaptation ability and our $\mathrm{Gather^{*}}$ indeed capture essential information in intermediate features.
@@ 4.2 Experiments on Fine-Grained Few-Shot Learning
To further assess the few-shot adaptation ability of MIET, following [16] we run MIET on fine-grained few-shot (FGFS) benchmark, which contains 5 fine-grained datasets: Aircraft [20], Food101 [2], Pets [23], Flowers102 [21] and Cars [17]. We fine-tune the edge network with train set containing 1, 2, 4, 8, 16-shot per class. Experiments are repeated three times with three different random seeds, and we report the average test set accuracy.
@@ Results.
As shown in Figure 4, the average few-shot accuracy of MIET consistently outperforms other 5 baseline methods in all cases. The average improvement of MIET compared to previous state-of-the-art FacT gets to 3.2% in 16-shot, which is the greatest improvement among all settings.
@@ 4.3 Experiments on Domain Generalization
Following the setting in [35], we test the robustness of MIET to domain shifts [36] when they are inevitable. We randomly sample 16 images from each class in ImageNet-1K to form train set. Apart from the validation set from ImageNet-1K [26], we also test the fine-tuned model on four out-of-domain datasets: 1) ImageNet-V2 [25] is collected from different sources than ImageNet but following the same collection protocol, 2) ImageNet-Sketch [31] is composed of sketch images of the same 1,000 classes in ImageNet, 3) ImageNet-A [12] contains adversarially-filtered images, 4) ImageNet-R [11] contains various artistic renditions of ImageNet-1K. We repeat the experiments for three times with different seeds and report average result over three seeds.
@@ Results.
As shown in Table 2, MIET outperforms previous methods consistently in both the source domain and in generalization to other domains. Compared to the previous the state-of-the-art method NOAH, the accuracy gain of MIET is very significant in ImageNet, ImageNet-Sketch,ImagnetNet-A and ImageNet-R, which are 5.0, 13.1, 9.5 and 9.0 percentage points, respectively. These result verifies the robustness of MIET against test set domain shift.
Table 3: Ablations on removing important components of MIET. “Avg.” denotes the group wise average accuracy on VTAB. $\mathrm{Gather}^{*}$ for (add intermediate features together / use output feature alone). LAE for (use low-rank attention edge network / linear probing).
@@
Table 4: Ablations on different block functions with ViT-B/16 as the pretrained model. # Layers denotes the number of layers in edge network. Avg. denotes the group wise average accuracy on VTAB.
@@ 4.4 Ablation studies
Now we conduct ablation studies on different components of MIET.
@@
we test the effectiveness of minimal interaction feature gathering ($\mathrm{Gather}^{*}$) and low-rank edge network in our method. Results are shown in Table 4.3. When both of them are not adopted, the method degrade to naive linear probing. When only minimal interaction feature gathering is used, then intermediate features are added for linear probing, the result in the second row reveal a 1.7% gain over naive linear probing, but this improvement is limited. When only LAE is used, the accuracy gain over naive linear probing is 9.1%, which indicates that adding an edge network directly after a pretrained model is useful for visual adaptation, but still lags far behind other fine-tuning methods. When both $\mathrm{Gather}^{*}$ and LAE are adopted, the average accuracy increases to 76.7% (on par with or better than other fine-tuning methods). This ablation shows that both components are important for MIET in visual adaptation, and only when two components are used together can we get the best results.
@@
We substitute the block function of edge network with MLP and Depthwise convolution (DWConv) [5]. Results are shown in Table 4.3. We can see that while low-rank attention (LAE) and MLP have the same number of parameters, LAE outperforms MLP with about 1% average accuracy lead. When the block function is changed to depthwise convolution, the number of parameters decreases, but the average accuracy also drops by a large margin. Overall, low-rank attention leads to the best performance and parameter efficiency tradeoff.
@@
To verify the effect of domain migration with different $k$, we test the respective accuracy on all 19 VTAB datasets for $k=6$, $k=9$ and $k=12$. Results are shown in Figure 5. For Natural and Specialized tasks, $k=12$ (adding all intermediate features) usually lead to the best result. But for structured datasets which have more low-level features, $k=9$ outperforms $k=12$ in 6 out of 9 datasets. These results indicate that in datasets with rich low-level features, intermediate outputs from early layers of the pretrained model are more important.
@@ 5 Related Work
Another line for visual adaptation is about memory-efficiency. LST [27] proposes to fine-tune a side network which is disentangled from pretrained model, and save memory by freeing pretrained model from back-propagation. DTL [7] considers adding side information back to pretrained network to make a tradeoff between adaptation ability and memory consumption. LAST [29] proposes a low-rank attention side network for effective visual adaptation.
6 Conclusion and Discussion
In this paper, we proposed a visual adaptation paradigm and a method which allows edge devices with extremely low computation resources to finetune large vision model without access to the pretrained model itself. While a pretrained model resides in cloud and provides features that are too large to transfer to the edge device, our minimal interaction edge tuning (MIET) compresses the features while keeping essential information for downstream tasks. The entire fine-tuning process is performed only on the edge device, requesting only limited storage, compute, memory, and communication resources while achieving high accuracy. In particular, we emphasized the communication factor, which has been largely ignored in previous research.
Currently, our MIET is restricted to visual recognition tasks. However, side-tuning has been shown to be able to work for object detection. We will work on this limitation, by extending MIET to object detection and other visual tasks. More vision tasks, such as image and video generation, may also enjoy benefits brought up by MIET. We will also explore other gather functions in the minimal interaction side tuning framework, which may provide even higher accuracy and smaller interaction.
Appendix A Detailed results on VTAB-1K
@@
MIET 67.8 93.2 76.8 99.7 91.1 87.5 54.0 87.3 95.8 87.1 75.3 84.5 66.8 51.2 83.7 84.0 45.7 33.3 50.3 76.7
Appendix B Imlementation details
@@
Appendix C Experiment on varying side input for side tuning

### v3

@@ 4.1 Experiments on VTAB-1K
We compare MIST with various fine-tuning methods. Basically, we divide them into two categories: those potentially support separated tuning and those do not. For the first category, we have linear probing, Partial-1 (which unfreeze the final layer of ViT-B), LST [27], LAST [29] and our method MIST. For the second category, we have BitFit [34], VPT [15], LoRA [14], AdaptFormer [4], FacT [16], SPT-LoRA [8].
@@ Results.
The results on VTAB-1K are shown in Table 1. Methods that potentially support separated tuning are marked with $\checkmark$ in corresponding cells. In the “Transfer Overhead” column, we list the communication cost per image between cloud and low-resource devices. MIST only requires to transfer 0.58MB of information per image, lower than other compared methods (except linear probing). This overhead is only about $1/7$ of LAST and $1/13$ of LST. Hence, MIST has met our “minimal interaction” goal.
With respect to trainable parameters, MIST only introduces 0.38M trainable parameters, lower than other methods that support separated tuning. Though MIST requires slightly more trainable parameters than some non-separated tuning methods, this difference is in fact negligible with respect to parameter size of the pretrained model (e.g., MIST: 0.38/85.8=0.44% vs. FacT: 0.07/85.8=0.08%, both $<1\%$).
We also list the MACs and GPU memory footprint. For non-separated tuning methods, the whole model are enforced to be fine-tuned on low-resource devices. For methods that support separated tuning, the pretrained model is deployed on a cloud server and side networks are deployed on low-resource devices, leading to very low MACs and memory footprint on low-resource devices. MIST requires the lowest MACs and GPU memory (again except linear probing) on low-resource devices, which are respectively about 1% and 10% of normal PEFT methods.
MIST achieves the highest group-wise accuracy compared to other methods. It has the highest group accuracy for specialized datasets and structured datasets. This result proves that MIST has strong adaptation ability and our $\mathrm{Gather^{*}}$ indeed captures essential information in intermediate features.
@@ 4.2 Experiments on Fine-Grained Few-Shot Learning
To further assess the few-shot adaptation ability of MIST, following [16] we ran MIST on the fine-grained few-shot (FGFS) benchmark, which contains 5 fine-grained datasets: Aircraft [20], Food101 [2], Pets [23], Flowers102 [21] and Cars [17]. We fine-tuned the edge network with train sets containing {1, 2, 4, 8, 16}-shots per class. Experiments were repeated three times with three different random seeds, and we report the average test set accuracy.
@@ Results.
As shown in Figure 4, the average few-shot accuracy of MIST consistently outperforms other 5 baseline methods in all cases. Average improvement of MIST compared to previous state-of-the-art, FacT, reaches 3.2% in 16-shot, which is the greatest improvement among all settings.
@@ 4.3 Experiments on Domain Generalization
Following the setting in [36], we tested the robustness of MIST to domain shifts [37] when they were inevitable. We randomly sampled 16 images from each class in ImageNet-1K to form the train set. Apart from the validation set from ImageNet-1K [26], we also tested the fine-tuned model on four out-of-domain datasets: 1) ImageNet-V2 [25] which is collected from different sources than ImageNet but following the same collection protocol, 2) ImageNet-Sketch [31] which is composed of sketch images of the same 1,000 classes in ImageNet, 3) ImageNet-A [12] which contains adversarially-filtered images, 4) ImageNet-R [11] which contains various artistic renditions of ImageNet-1K. We repeated the experiments for three times with different seeds and report average result over three seeds.
@@ Results.
As shown in Table 2, MIST outperforms previous methods consistently in both the source domain and in generalization to other domains. Compared to the previous state-of-the-art method NOAH, the accuracy gain of MIST is significant in all ImageNet, ImageNet-Sketch, ImageNet-A and ImageNet-R, which are 5.0, 13.1, 9.5 and 9.0 percentage points, respectively. These result verifies the robustness of MIST against test set domain shift.
@@
Table 3: Ablations on removing important components of MIST. “Avg.” denotes the group wise average accuracy on VTAB. $\mathrm{Gather}^{*}$ for adding intermediate features together or using output feature alone. LAE for using our low-rank attention edge network or linear probing).
@@
Table 4: Ablations on different block functions with ViT-B/16 as the pretrained model. #Layers denotes number of layers in edge network. Avg. denotes group-wise average accuracy on VTAB.
4.4 Experiments on Edge Devices
To validate the capability of our method on edge devices, we conducted the fine-tuning and inference of the ViT-B model on a Raspberry Pi 4B board with only 2GB SDRAM (with neither NPU nor GPU). During the fine-tuning process, with a batch size of 32 and an image size of 224x224, the edge device required only 663 MB peak memory usage in the whole finetuning process. During inference, the program on the edge device required only 255 MB of memory. This level of resource consumption allows the algorithm to run smoothly on many small embedded devices.
The experiment on real world edge devices (in our case a Raspberry Pi) effectively demonstrates the capability of our method to update models in real-time on edge devices and other low-resource devices.
@@ 4.5 Ablation studies
We also conducted ablation studies on different components of MIST.
@@
we tested the effectiveness of minimal interaction feature gathering ($\mathrm{Gather}^{*}$) and low-rank edge network in our method. Results are shown in Table 3. When both are not adopted, the method degrade to naive linear probing. When only minimal interaction feature gathering is used, then intermediate features are added for linear probing. The result in the second row shows a 1.7% gain over naive linear probing, but this improvement is limited. When only LAE is used, the accuracy gain over naive linear probing is 9.1%, which indicates that adding an edge network directly after a pretrained model is useful for visual adaptation, but still lags far behind other PEFT methods. When both $\mathrm{Gather}^{*}$ and LAE are adopted, the average accuracy increases to 76.7% (on par with or better than other fine-tuning methods). This ablation shows that both components are important for MIST in visual adaptation, and only when two components are used together can we get the best results.
@@
We substituted our LAE block function in the edge network with MLP or Depthwise convolution (DWConv) [5]. Results are shown in Table 4. We can see that while low-rank attention (LAE) and MLP have the same number of parameters, LAE outperforms MLP with about 1% average accuracy lead. When the block function is changed to depthwise convolution, the number of parameters decreases, but the average accuracy also drops by a large margin. Overall, low-rank attention leads to the best performance and parameter efficiency tradeoff.
@@
To verify the effect of domain migration with different $k$, we tested the respective accuracy on all 19 VTAB datasets for $k=6$, $k=9$ and $k=12$. Results are shown in Figure 5. For Natural and Specialized tasks, $k=12$ (adding all intermediate features) usually lead to the best result. But for ‘Structured’ which have more low-level features, $k=9$ outperforms $k=12$ in 6 out of 9 datasets. These results indicate that in datasets with rich low-level features, intermediate outputs from early layers of the pretrained model are more important.
@@ 5 Related Work
Another line for visual adaptation is about memory-efficiency. LST [27] proposes to fine-tune a side network which is disentangled from pretrained model, and save memory by freeing pretrained model from back-propagation. DTL [7] considers adding side information back to pretrained network to make a tradeoff between adaptation ability and memory consumption. LAST [29] proposes a low-rank attention side network for effective visual adaptation. MIST stems from this line of work, but elaborates on minimizing the interaction between backbone and side networks, which previous side tuning methods (like LAST) were not capable of, but is the key to realizing the proposed ‘Separated Tuning’.
Another related work is Offsite-Tuning [32]. Offsite-Tuning extracts a small version of the base model called ‘emulator’ to emulate the behavior of the base model, and fine-tune adapter with the emulator rather than the base model. This helps to reduce computation overhead during fine-tuning and protect data privacy.
6 Conclusions and Discussions
In this paper, we proposed a visual adaptation paradigm and a method which allows devices with low computational resources to finetune large vision model without access to the pretrained model itself. While a pretrained model resides in cloud and provides features that are too large to transfer to the low-resource device, our minimal interaction separated tuning (MIST) compresses the features while keeping essential information for downstream tasks. The entire fine-tuning process is performed only on the low-resource device, requesting only limited storage, compute, memory, and communication resources while achieving high accuracy. In particular, we emphasized the communication factor, which has been largely ignored in previous research.
Currently, experiments on MIST are mostly restricted to visual recognition tasks, and this might constrain the scope of usage of MIST. However, we have adapted MIST to segmentation. And the results showed that MIST is also suitable for higher level visual tasks (cf. appendix for results). More vision tasks, such as image and video generation, may also enjoy benefits brought up by MIST. We will also explore other gather functions in the new minimal interaction separated tuning framework, which may provide even better overall performance.
Supplementary Material
7 Detailed results on VTAB-1K
@@
8 Segmentation result on ADE-20k
For semantic segmentation, we fine-tune a SegFormer-B4 model on ADE20K dataset. SegFormer has a multi-stage architecture, where the backbone is divided into 4 stages. To ensure compatibility for feature addition, we use the downsampling module in each stage to downsample the spatial dimensions of the aggregated features, but without training. This aggregated feature is used as the final feature from backbone and fine-tuned with LAE. Due to the computational limitation, we only fine-tune 42k iters for all methods, which is roughly 1/4 of the 160k iters mentioned in segformer paper. MIST achieves 2.4 mIoU gain compared to fine-tuning with frozen encoder (42.9 vs. 40.5), and is only 0.5 point behind full fine-tuning (42.9 vs. 43.4). This result shows that MIST also work
Table 5: Ablations of semantic segmentation. On ADE20K dataset with SegFormer-b4. We use mIoU as evaluation metric. Params denotes number of trainable parameters in SegFormer encoder. Gather for (add intermediate features together / use output feature alone). LAE for (use low-rank attention edge network / not use).
Method Gather LAE Params. mIoU
Frozen - - 00.0M 40.5
Full-FT - - 60.8M 43.4
MIST 00.0M 40.7
MIST 00.7M 42.3
MIST 00.7M 42.9
@@ 8 Segmentation result on ADE-20k
MIST 67.8 93.2 76.8 99.7 91.1 87.5 54.0 87.3 95.8 87.1 75.3 84.5 66.8 51.2 83.7 84.0 45.7 33.3 50.3 76.7
9 Imlementation details
@@
10 Experiment on varying side input for side tuning